Sally has a piece of cloth. She used 1/3 of the cloth to make a pillow cover and another 4/7 of the cloth to make a tablecloth.
a. What fraction of the cloth was used altogether for the pillow cover and tablecloth?

Answer by stanbon(75887)

Sally has a piece of cloth. She used 1/3 of the cloth to make a pillow cover and another 4/7 of the cloth to make a tablecloth.
a. What fraction of the cloth was used altogether for the pillow cover and tablecloth?
----------------------
Ans:: (1/3) + (4/7) = (7 + 12)/21 = 18/21
------
Cheers,
Stan H.
-----------